Madkpow:

For an experiment, Alicia plans to spin two different spinners, with equal-sized sections. The first spinner has sections labeled red, orange, yellow, green, and blue. The second spinner has sections labeled with the numbers 1 to 5. 1 2 3 4 5 Red Orange Yellow Green Blue Use the area model above to determine the probability that the first spinner will land on orange or green and the second spinner will land on an even number.

Aries:

Okay, so we know that the first spinner must land on orange or green and the second one an even number. So out of 1-5, there are two even numbers, 2 and 4. So we have two numbers and two colors. If the first spinner land on orange, then the second one can land on 2 or 4, same with green. So we get four probabilities. Green 2, Green 4, Orange 2, and Orange 4.

K1NGofPadlet:

Let's see. There are five sections on the first spinner and two of them are orange and green, so the probability of the first spinner landing on orange or green is 2/5. The second spinner has five sections labeled 1 to 5, and two of the numbers are even, 2 and 4. So, the probability of the second spinner landing on an even number is 2/5. To find the probability that both events will occur, we can multiply the two probabilities together: Probability of landing on orange or green = 2/5 Probability of landing on an even number = 2/5 Probability of both events occurring = (2/5) x (2/5) = 4/25 Therefore, the probability that the first spinner will land on orange or green and the second spinner will land on an even number is 4/25.

KyledaGreat:

To use the area model, we can list all the possible outcomes in a grid and shade the outcomes that satisfy the conditions given: 1 2 3 4 5 Red Orange ■ ■ Yellow Green ■ ■ Blue In the grid above, we have shaded the outcomes where the first spinner lands on orange or green and the second spinner lands on an even number. There are a total of 10 possible outcomes, and 4 of them satisfy the conditions we are interested in. Therefore, the probability that the first spinner will land on orange or green and the second spinner will land on an even number is: P(orange or green and even) = 4/10 = 2/5 = 0.4 So the probability is 0.4 or 40%.

Renne:

Alicia is conducting an experiment where she will be spinning two different spinners with equal-sized sections. The first spinner has sections labeled with colors: red, orange, yellow, green, and blue. The second spinner has sections labeled with numbers from 1 to 5. To calculate the probability of the first spinner landing on orange or green and the second spinner landing on an even number, we can use the area model shown above. By counting the total number of possible outcomes, we see there are 25 total sections. Out of those sections, there are 10 sections where the first spinner lands on orange or green, and there are 10 sections where the second spinner lands on an even number. There are 4 sections where both events occur, which means the probability of both events occurring is 4/25 or 16%.